A publishing company claims that in fall 2019, the average price of college textbooks for a single semester is $385. Suppose you decide to collect data from a random sample of students to assess whether the publisher's claim is reasonable, and you find that in a random sample of 22 college students, the mean price of textbooks for the fall 2019 semester was $433.50 with a standard deviation of $86.92.  

At the 0.01 significance level, is there sufficient evidence to conclude that the mean price of college textbooks for a single semester is different from the value claimed by the publisher?  We can assume that the population of textbook prices are normally distributed.

What is the standardized test statistic? Round to two decimal places.
Answer #1

Here we have: x = 433.50 S = 86.92 n = 22 a = 0.01 Hypothesis: Ho; u = 385 Ha; + 385 Test statistics: t-distribution (o unkno

tarit= ta/21df = 10.01/2,21 = +2.831 P-value = 2p (t>|2.621) = 0.0160 Fail to reject the null hypothesis. Conclusion: Here we